Ama's age is two - thirds that of her elder sister. Four years ago ama's age was half the sisters.how old is the sister
Answer by fractalier(6550)

Call Ama's age NOW, a. Her sister's age NOW will be s. Then we have
a = (2/3)s
a - 4 = (1/2)(s - 4)
Substitute the first into the second equation and get
(2/3)s - 4 = (1/2)(s - 4)
Multiply by 6 to clear fractions and get
4s - 24 = 3(s - 4)
4s - 24 = 3s - 12
Subtract 3s and add 24 and get
s = 12
a = (2/3)(12) = 8
Ama is 8, the sister is 12.